Help pls. It's confusing.
Karla weighs 42 lbs more than mara. Their combined weight is 68lbs. less than three times Mara's weight. How much does Karla weigh?

Answer by rfer(16322)

mara=x=110
karla=x+42=152
total=3x-68
--------------
x+(x+42)=3x-68
2x+42=3x-68
-x=-110
x=110